OpenStudy (anonymous):

The perimeter of a rectangular sign is not to exceed 50ft . The length is to be twice the width. What widths will meet these conditions?

OpenStudy (anonymous):

L=2W P=2(L+W) P=2(2W+W) P=6W=50 W=50/6 L=50/3